OpenStudy (anonymous):

Which of the following is equivalent to 5^-3

OpenStudy (campbell_st):

its really \[\frac{1}{5^3} ... or..... \frac{1}{125}\] negative powers indicate fractions.

OpenStudy (anonymous):

those aren't a choice @campbell_st

OpenStudy (anonymous):

are you giving an online test? @jayfer

OpenStudy (anonymous):

no @gsRttN homework and i dont understand these kinda questions

OpenStudy (dls):

post the options

OpenStudy (anonymous):

@jayfer post the choices

OpenStudy (unklerhaukus):

\[5^{-3}=\frac15\times\frac15\times\frac15\]

OpenStudy (skullpatrol):

0.008